The ratio of boys to girls at an elementary school is 3 to 4. If there are 580 students, how many girls are in the school?
Answer by ankor@dixie-net.com(22740)

The ratio of boys to girls at an elementary school is 3 to 4.
If there are 580 students, how many girls are in the school?
:
This total (580) for students can not have a 3:4 ratio
I will assume a value of 581 students to make the problem work
:
Let x = the multiplier
then
3x = no. of boys
4x = no. or girls
:
3x + 4x = 581
7x = 581
x = 581/7
x = 83 is the multiplier
then
4*83 = 332 girls
:
:
Check this by finding the number of boys; 3*83 = 249 boys
249 + 332 = 581